The atmosphere in a sealed diving bell contained oxygen and helium. If the gas mixture has 0.170 atm of oxygen and a total pressure of 3.20 atm, calvulate the mass of helium in 10.0 L of the gas mixture at 23 degrees C.

Mass= g
Answer #1

Total pressure of O2 and He(PT) = 3.20atm

The partial pressure of O2      = 0.17atm

partial pressure (PHe)    = PT-PO2

                                        = 3.20-0.17

                                         = 3.03atm

PHe   = 3.03atm

V = 10L

T   = 23+273   = 296K

PV = nRT

n   = PV/RT

      = 3.03*10/(0.0821*296)

       = 1.25moles

mass of He = no of moles * gram atomic mass

                   = 1.25*4

                    = 5g >>>.answer